OpenStudy (anonymous):

A car is traveling at a constant velocity. After two hours the car has traveled 90 miles. After 5 hours the car has traveled 225 miles. Assuming a linear equation fits this situation, how far will the car have traveled after seven hours?

OpenStudy (lgbasallote):

\[\large speed = \frac{distance}{time} = \frac{90}{2} = 45 mph\] \[\large distance = speed \times time = 45 \times 7\]

OpenStudy (anonymous):

So the answer is 315?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Thanks(: